# Learning to Partially Defer for Sequences

### v1

@@ Abstract
In the Learning to Defer (L2D) framework, a prediction model can either make a prediction or defer it to an expert, as determined by a rejector. Current L2D methods train the rejector to decide whether to reject the entire prediction, which is not desirable when the model predicts long sequences. We present an L2D setting for sequence outputs where the system can defer specific outputs of the whole model prediction to an expert in an effort to interleave the expert and machine throughout the prediction. We propose two types of model-based post-hoc rejectors for pre-trained predictors: a token-level rejector, which defers specific token predictions to experts with next token prediction capabilities, and a one-time rejector for experts without such abilities, which defers the remaining sequence from a specific point onward. In the experiments, we also empirically demonstrate that such granular deferrals achieve better cost-accuracy tradeoffs than whole deferrals on Traveling salesman solvers and News summarization models.
@@ 1 Introduction
Trustworthiness of AI systems is under increased scrutiny with the increased deployment of such systems in highly consequential areas like healthcare [1, 2, 3] and criminal justice [4, 5, 6]. Hybrid intelligent systems [7, 8, 9, 10] approach this problem through collaborations between either humans and machines or two machines to improve confidence in the system. Learning to Defer (L2D) [11] is one such framework that accommodates these teams by allowing the model to defer to an expert when it is uncertain about a prediction task with the objective of maximizing overall system accuracy while minimizing deferral costs.
Existing L2D methods focus on multiclass classification [12, 13, 14, 15, 16] and scalar regression [17, 18]. However, these methods only support complete deferral of prediction outputs, resulting in costs that scale with the output size, when applied to large sequence outputs [19]. This is especially inefficient if only parts of the model’s prediction are inaccurate. The prevalence of sequences in areas such as drug discovery [20, 21, 22, 23] and language modeling [24, 25, 26] calls for a cost-effective extension of the L2D framework to handle sequential outputs.
@@
While partial deferral is therefore desirable, the granularity of such deferrals depends on the size of the sequence segments the expert can predict at a time. For instance, if the expert is a next-token predictor, it can complete deferred parts of the prediction, as small as a single token. This form of partial deferral that we call token-level deferral can become especially useful when the model’s predictive quality across all parts of the output are dependent. Specifically, in autoregressive prediction, the uncertainty at earlier prediction points can propagate to later points and, in turn, affect the predictive accuracy of the complete sequence [27, 28, 29]. Thus, to achieve a better cost-accuracy tradeoff, the token-level rejector can orchestrate a dynamic collaboration between the predictor and the expert where the predictor receives an expert prediction on an early substructure to predict the next token before the model veers “off-course”.
Although the granularity offered from token-level deferral can improve predictive quality with low deferral costs, this is only feasible with experts that can make predictions one token at a time. That said, partial deferrals can still be implemented as long as the expert is at least capable of completing the predicted sequence given a partial subsequence from the model. We refer to this type of deferral as one-time deferral. For example, Gurobi solvers for Traveling salesman problems lack the ability to iteratively select the next city in the tour, but they can use partial subtours generated by models as constraints to generate the entire tour at a reduced cost. In this case, the one-time rejector would select a point in the sequence from which the expert completes the sequence.
@@
  • We present a novel learning to defer framework where the rejector interweaves the expert’s and machine’s predictions to improve cost-accuracy tradeoffs. We divide this setting into token-level and one-time deferrals based on the granularity of the deferrals, thereby accommodating various types of experts.
  • We propose convex surrogate loss functions that offer Bayes consistency guarantees leading to implementable model-based post hoc rejectors for pre-trained predictors for each type of deferral. We further analyze the surrogates theoretically through generalization bounds.
  • We empirically demonstrate improved cost-accuracy tradeoffs when using model-based rejectors for either deferrals against whole sequence rejectors with news summarization and traveling salesman problem solvers.
@@ 2.1 Learning to Defer
Deferrals, also known as rejections or abstentions, are classified into two types - confidence-based [30, 31, 32, 33] and model-based [34, 11, 12, 14, 13, 15, 17, 18].
Confidence-based methods involve learning a predictor where there is some prescribed cost for rejection; if the confidence score outputted by the learned predictor falls below a cost-dependent threshold, then it is rejected. However, Madras et al. [11] and Cortes et al. [34] characterized the suboptimality of these methods and advocated for the rejector to be a model trained either simultaneously with the predictor or post-hoc: this is the model-based approach. This system seeks to minimize the following loss function:
$\begin{split}{\cal L}_{\text{L2D}}&\left(h,r,e,x,y\right)\\ &=l(y,h(x))\mathbbm{1}_{r(x)<0}+c(x,y,e)\mathbbm{1}_{r(x)\geq 0}\end{split}$ (1)
where $h$ is the predictor, $r$ is the rejector, $e$ is the expert that the model defers to, $l(y,h(x))$ is the model loss, and $c(x,y,e)$ is the cost incurred when deferring to an expert.
While our settings assume access to only one expert, we note that one-time deferral closely resembles L2D with access to predictions from multiple experts [35, 18]. The details of this similarity and its implications are discussed in Section 4.1. In particular, L2D for $E$ experts aims to minimize the following loss function:
$\begin{split}&{\cal L}_{\text{L2DMulti}}\left(h,r,x,y,\{e_{j}\}_{j=1}^{E}\right)\\ &=l(y,h(x))\mathbbm{1}_{r(x)=0}+\sum_{j=1}^{E}c_{j}(x,y,e_{j})\mathbbm{1}_{r(x)=j}\end{split}$ (2)
where $c_{j}(x,y,e_{j})$ is the cost of querying the $j^{\text{th}}$ expert, $e_{j}$. The rejector in this scenario now must decide whether to defer and which expert to defer to.
@@ 2.2 Consistent Losses
The direct minimization of Equations 1 and 2 can be challenging due to the non-convex and discontinuous nature of the functions. This can be addressed by optimizing a convex surrogate loss function which, upon minimization, will lead to the same outcome as minimizing the original loss. This property is formally called Bayes Consistency.
@@
Let the risk of a hypothesis $f$ with respect to a loss $l$ be ${\cal R}_{l}(f)=\mathbb{E}_{(x,y)\sim{\cal P}_{{\cal X},{\cal Y}}}[l(f(x),y)]$. Let the optimal risk with respect to $l$ out of all hypotheses in ${\cal F}$ be ${\cal R}_{l}^{*}({\cal F})=\inf_{f\in{\cal F}}{\cal R}_{l}(f)$. A surrogate loss function, $\phi(\cdot)$, is said to be ${\cal F}$-consistent with respect to $l$ if for any $f_{n}\in{\cal F}$, the following property holds true:
@@
When ${\cal F}$ is the class of all measurable functions, Bayes consistency is achieved if the excess risk of $l$ can be related to the excess risk of $\phi$ through a non-decreasing function $\Gamma:\mathbb{R}\rightarrow\mathbb{R}$ for which $\Gamma$ is continuous at zero and $\Gamma(0)=0$ in the following manner:
@@
where ${\cal R}_{l}^{*}=\inf_{f}{\cal R}_{l}(f)$ is the optimal risk over the class of all measurable functions. In this paper, we propose loss functions similar to Equations 1 and 2 for both token-level and one-time deferral and then, we show that the above inequality holds true for our surrogate designs for both forms of deferral, proving Bayes consistency.
@@ 3 Token-level partial deferral
Let ${\cal Y}\subseteq{\cal V}^{L}$ be the output sequence space where each sequence can have a maximum length $L$ and ${\cal V}$ is the label set that any individual component or token $y_{j}\in y\in{\cal Y}$ of a sequence can belong to. ${\cal V}$ can be a finite vocabulary of tokens in the case of text sequences and be $\mathbb{R}$ when the output is a scalar time series. Let ${\cal X}$ be the feature space. In sequence generation, ${\cal X}\subseteq\mathbb{R}^{d}$ would be a $d$-dimensional feature space, whereas ${\cal X}={\cal V}^{\prime L}$, perhaps, is characterized by a distinct ${\cal V}^{\prime}$ vocabulary set in sequence-to-sequence learning tasks. We denote ${\cal P}_{X,Y}$ to be the data-generating distribution over ${\cal X}\times{\cal Y}$.
@@
Suppose we have a predictor $h$ and an expert $e$ with stronger predictive capabilities than $h$. $e$ takes in $x\in{\cal X}$ and the leftward context $\widehat{y}_{<j}=(\widehat{y}_{1},\cdots,\widehat{y}_{j-1})\in{\cal V}^{j-1}$ as inputs to predict the $j^{\text{th}}$ label in the sequence; repeated calls to $e$ can auto-regressively complete the entire sequence prediction task. $e$ can, for instance, be a human or a large autoregressive model. However, querying $e$ is more expensive than calling $h$. As a result, instead of deferring an entire sequence prediction to the expert, we wish to design an L2D system where experts can complete uncertain tokens in the model’s prediction, facilitating more fine-grained deferrals - referred to as token-level deferrals - in order to achieve a better cost-quality tradeoff.
@@
If $h$ can also perform autoregressive prediction, token-level deferrals can additionally enable the predictor to leverage the predicted token from the expert to predict the next token, fostering a more seamless and adaptive collaboration between the expert and the predictor. This can further improve cost-effectiveness by curbing uncertainty propagation for the predictor, thereby reducing the number of expert calls. While our token-level deferrals paradigm can apply to all sequence predictors, we will specifically focus on creating a dynamic L2D system with an autoregressive predictor.
@@
In this setting, at the $j^{\text{th}}$ step, the rejector for the $j^{\text{th}}$ token, $r_{j}$, takes in $x$ and $\widehat{y}_{<j}$. It is important to note that $\widehat{y}_{<j}$ can be viewed as a mixture of predictions from the model and the expert. When $r_{j}(x,\widehat{y}_{<j})\geq 0$, the rejector chooses to reject the $j^{\text{th}}$ label prediction from $h$ and receives a prediction $\widehat{y}_{j}^{e}$ from an expert. Otherwise, the model will continue with the prediction from $h$, i.e. $\widehat{y}_{j}^{h}=h(x,\widehat{y}_{<j})$. For a fixed $h$ and a specific instance of $(x,\widehat{y}_{<j},y)$, the optimal $r_{j}$ can be determined by minimizing the following loss function:
$\begin{split}{\cal L}_{j}\left(h,r_{j},x,\widehat{y}_{<j},y\right)=l&\left(y,\widehat{y}_{j}^{h}\right)\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)<0}\\ &+c_{j}(x,\widehat{y}_{<j},y)\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)\geq 0}\end{split}$ (3)
@@
where $l\left(y,\widehat{y}_{j}^{h}\right)$ is the loss incurred from using $\widehat{y}_{j}^{h}$ and $c_{j}(x,\widehat{y}_{<j},y)$ is the cost of querying the expert for the $j^{\text{th}}$ token. $l\left(y,\widehat{y}_{j}^{h}\right)$ can be the mean squared error defined by $(y_{j}-\widehat{y}_{j}^{h})^{2}$ or a 00-$1$ loss between the true $j^{\text{th}}$ token and $\widehat{y}_{j}^{h}$ in text sequences. Generally, $l\left(y,\widehat{y}_{j}^{h}\right)$ can be viewed as token-level feedback to generate an accurate complete sequence. $c_{j}(x,\widehat{y}_{<j},y)$ can be a constant $c_{j}$ to reflect the cost of querying $e$ at the $j^{\text{th}}$ step but it can also be designed as $l\left(y,\widehat{y}_{j}^{e}\right)$ to account for the expert’s errors.
@@
Finally, the predicted $j^{\text{th}}$ label $\widehat{y}_{j}=\widehat{y}_{j}^{h}\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)<0}+\widehat{y}_{j}^{e}\mathbbm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)\geq 0}$ is then appended to $\widehat{y}_{<j}$ to serve as an input for $(j+1)^{\text{th}}$ token prediction, and this process continues until $L$ labels in the sequence are generated. The total deferral loss for the entire sequence can be characterized as ${\cal L}^{\text{Token}}\left(h,r,x,y\right)=\frac{1}{L}\sum_{j=1}^{L}{\cal L}_{j}\left(h,r_{j},x,\widehat{y}_{<j},y\right)$ where $r=\begin{bmatrix}r_{1}&\dots&r_{L}\end{bmatrix}^{T}$. The ${\cal L}^{\text{Token}}$-risk is denoted by ${\cal R}_{\text{Token}}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\text{Token}}(h,r,x,y)]$.

### v2

@@ Abstract
In the Learning to Defer (L2D) framework, a prediction model can either make a prediction or defer it to an expert, as determined by a rejector. Current L2D methods train the rejector to decide whether to reject the entire prediction, which is not desirable when the model predicts long sequences. We present an L2D setting for sequence outputs where the system can defer specific outputs of the whole model prediction to an expert in an effort to interleave the expert and machine throughout the prediction. We propose two types of model-based post-hoc rejectors for pre-trained predictors: a token-level rejector, which defers specific token predictions to experts with next token prediction capabilities, and a one-time rejector for experts without such abilities, which defers the remaining sequence from a specific point onward. In the experiments, we also empirically demonstrate that such granular deferrals achieve better cost-accuracy tradeoffs than whole deferrals on Traveling salesman solvers, News summarization, and Weather prediction.
@@ 1 Introduction
Trustworthiness of AI is under scrutiny with its increased deployment in highly consequential areas like healthcare (Ker et al., 2017; Courtiol et al., 2019; Asan et al., 2020) and criminal justice (Dressel and Farid, 2018; Rigano, 2019; Alikhademi et al., 2022). Hybrid intelligent systems (Kamar, 2016; Dellermann et al., 2019; Akata et al., 2020; Maadi et al., 2021) approach this problem through collaborations between either humans and machines or two machines to improve confidence in the system. Specifically, Learning to Defer (L2D) (Madras et al., 2018) accommodates these teams by allowing the model to defer to an expert when it is uncertain about a prediction task with the objective of maximizing overall system accuracy while minimizing deferral costs.
Existing L2D methods focus on multiclass (Mozannar and Sontag, 2020; Cao et al., 2022; Verma and Nalisnick, 2022; Mao et al., 2024a, c) and scalar regression (Cheng et al., 2024; Mao et al., 2024b) problems. However, these methods only support complete deferral of predictions, resulting in costs that scale with the output size, when applied to large sequence outputs (Narasimhan et al., 2024). This is inefficient if only parts of the model’s prediction are inaccurate. The prevalence of sequences in areas such as drug discovery (Wang et al., 2018; Jumper et al., 2021; Dauparas et al., 2022; Nijkamp et al., 2023) and language modeling (Kenton and Toutanova, 2019; Brown et al., 2020; Raffel et al., 2020) calls for a cost-effective extension of the L2D framework to handle sequential outputs.
@@
While partial deferral is therefore desirable, the granularity of such deferrals depends on the size of the sequence segments the expert can predict at a time. For instance, if the expert is a next-token predictor, it can complete spans as small as a single token—token-level deferral. This is especially valuable when the model’s predictive quality across all parts of the output are dependent. Specifically, in autoregressive prediction, the uncertainty at earlier prediction points can propagate and degrade later points (Bengio et al., 2015; Lamb et al., 2016; Schmidt, 2019). Thus, to achieve a better cost-accuracy tradeoff, the token-level rejector can orchestrate a dynamic collaboration between the predictor and the expert where the predictor receives an expert prediction on an early substructure to predict the next token before the model veers “off-course”.
This, however, is only feasible with experts that can make predictions one token at a time. That said, partial deferrals can still be implemented as long as the expert is at least capable of completing the predicted sequence given a partial subsequence from the model. We refer to this type of deferral as one-time deferral. For example, Gurobi solvers for Traveling salesman problems lack the ability to iteratively select the next city in the tour, but they can use partial subtours generated by models as constraints to generate the entire tour at a reduced cost. In this case, the one-time rejector would select a point in the sequence from which the expert completes the sequence.
@@
  • We present a novel L2D framework where the rejector interweaves the expert’s and machine’s predictions to improve cost-accuracy tradeoffs. We divide this setting into token-level and one-time deferrals based on the granularity of the deferrals, thereby accommodating various types of experts.
  • We propose convex surrogate losses that offer Bayes consistency guarantees leading to implementable model-based post hoc rejectors for pre-trained predictors for each type of deferral. We further analyze the surrogates theoretically through generalization bounds.
  • We empirically demonstrate improved cost-accuracy tradeoffs when using model-based rejectors for either deferrals against whole sequence rejectors with Traveling salesman solvers, News summarization, and Weather prediction.
@@ 2.1 Learning to Defer
Deferrals, also known as rejections or abstentions, are classified into two types - confidence-based (Chow, 1957; Bartlett and Wegkamp, 2008; Yuan and Wegkamp, 2010; Ramaswamy et al., 2018) and model-based (Cortes et al., 2016; Madras et al., 2018; Mozannar and Sontag, 2020; Verma and Nalisnick, 2022; Cao et al., 2022; Mao et al., 2024a; Cheng et al., 2024; Mao et al., 2024b). Confidence-based methods learn a predictor that outputs confidences; a prediction is rejected whenever its confidence falls below a cost-dependent threshold. However, Madras et al. (2018) and Cortes et al. (2016) characterized the suboptimality of these methods and advocated for the rejector to be a model trained either simultaneously with the predictor or post-hoc: this is the model-based approach. This system seeks to minimize the following loss function:
${\cal L}_{\text{L2D}}\left(h,r,e,x,y\right)=l(y,h(x))\bm{1}_{r(x)<0}+c(x,y,e)\bm{1}_{r(x)\geq 0}$ (1)
where $h$ is the predictor, $r$ is the rejector, $e$ is the expert, $l(y,h(x))$ is the model loss, and $c(x,y,e)$ is the cost incurred when deferring to an expert.
Although our setting assumes access to a single expert, one-time deferral closely resembles L2D with multiple experts (Verma et al., 2023; Mao et al., 2024b). We formalize this similarity and its implications in Section 4.1. For $E$ experts, L2D minimizes:
${\cal L}_{\text{L2DMulti}}\left(h,r,x,y,\{e_{j}\}_{j=1}^{E}\right)=l(y,h(x))\bm{1}_{r(x)=0}+\sum_{j=1}^{E}c_{j}(x,y,e_{j})\bm{1}_{r(x)=j}$ (2)
where $c_{j}(x,y,e_{j})$ is the cost of querying the $j^{\text{th}}$ expert, $e_{j}$. The rejector must decide which expert to defer to.
@@ 2.2 Consistent Losses
Direct minimization of Equations 1 and 2 is challenging due to the non-convexity and discontinuities. This can be addressed by optimizing a convex surrogate loss function which, upon minimization, will lead to the same outcome as minimizing the original loss. This property is formally called Bayes Consistency.
@@
[…]
@@
[…]
@@
where ${\cal R}_{l}^{*}=\inf_{f}{\cal R}_{l}(f)$ is the optimal risk over all measurable functions. We introduce surrogates for both token-level and one-time deferral (similar to Equations 1 and 2) and prove they satisfy the above inequality, thereby establishing Bayes consistency.
@@ 3 Token-level partial deferral
Let $\mathcal{Y}\subseteq\bigcup_{\ell=1}^{L}\mathcal{V}^{\ell}$ denote the space of sequences over label set $\mathcal{V}$ with maximum length $L$. For $y\in\mathcal{Y}$, the $j^{\text{th}}$ token is $y_{j}\in\mathcal{V}$. $\cal V$} can be a finite vocabulary in the case of text sequences and be $\mathbb{R}$ when the output is a scalar time series. Let $\cal X$} be the feature space. In sequence generation, ${\cal X}\subseteq\mathbb{R}^{d}$ would be a $d$-dimensional feature space, whereas ${\cal X}={\cal V}^{\prime L^{\prime}}$ would be a sequence space in sequence-to-sequence learning. Let ${\cal P}_{X,Y}$ denote the data-generating distribution over ${\cal X}\times{\cal Y}$.
@@
Suppose we have a predictor $h$ and a stronger expert $e$. $e$ takes in $x\in{\cal X}$ and the leftward context $\widehat{y}_{<j}=(\widehat{y}_{1},\cdots,\widehat{y}_{j-1})\in{\cal V}^{j-1}$ to predict $y_{j}$; repeated calls can auto-regressively complete the sequence. $e$ can, for instance, be a human or a large autoregressive model. However, querying $e$ is more expensive than calls to $h$. We, therefore, wish to design an L2D system where experts can complete uncertain tokens in the model’s prediction, facilitating more fine-grained deferrals – referred to as token-level deferrals – yielding a more favorable cost-quality tradeoff than deferring the entire sequence.
@@
If $h$ is autoregressive, token-level deferrals allows $h$ to leverage the expert-supplied tokens when generating subsequent tokens, enabling a seamless and adaptive collaboration. This mitigates uncertainty propagation, reduces the number of expert queries, and improves cost–effectiveness. While this paradigm can apply to all sequence predictors, we will specifically focus on a dynamic L2D system with an autoregressive predictor.
@@
In this setting, at step $j$, the rejector $r_{j}$ takes in $(x,\widehat{y}_{<j})$ where $\widehat{y}_{<j}$ may mix predictor and expert tokens. When $r_{j}(x,\widehat{y}_{<j})\geq 0$, it rejects the $j^{\text{th}}$ label prediction from $h$ and receives an expert prediction $\widehat{y}_{j}^{e}$. Otherwise, the model will keep the prediction from $h$, i.e. $\widehat{y}_{j}^{h}=h(x,\widehat{y}_{<j})$. For a fixed $h$ and a specific instance of $(x,\widehat{y}_{<j},y)$, $r_{j}$ is trained to minimize:
${\cal L}_{j}\left(h,r_{j},x,\widehat{y}_{<j},y\right)=l\left(y,\widehat{y}_{j}^{h}\right)\bm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)<0}+c_{j}(x,\widehat{y}_{<j},y)\bm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)\geq 0}$ (3)
@@
where $l(y,\widehat{y}_{j}^{h})$ is the loss from using $\widehat{y}_{j}^{h}$ and $c_{j}(x,\widehat{y}_{<j},y)$ is the deferral cost. $l(y,\widehat{y}_{j}^{h})$ can be the mean squared error defined by $(y_{j}-\widehat{y}_{j}^{h})^{2}$ or a 0-11 loss between the true $j^{\text{th}}$ token and $\widehat{y}_{j}^{h}$ in text sequences. Generally, $l\left(y,\widehat{y}_{j}^{h}\right)$ can be viewed as token-level feedback to generate an accurate complete sequence. $c_{j}(x,\widehat{y}_{<j},y)$ can be a constant $c_{j}$ to reflect expert query costs but it can also be $l(y,\widehat{y}_{j}^{e})$ to account for the expert’s errors.
@@
Finally, the predicted $j^{\text{th}}$ label $\widehat{y}_{j}=\widehat{y}_{j}^{h}\bm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)<0}+\widehat{y}_{j}^{e}\bm{1}_{r_{j}\left(x,\widehat{y}_{<j}\right)\geq 0}$ is appended to $\widehat{y}_{<j}$ to serve as input for $(j+1)^{\text{th}}$ token prediction, and this continues until $L$ labels are generated. The total loss for the sequence is ${\cal L}^{\text{Token}}\left(h,r,x,y\right)=\frac{1}{L}\sum_{j=1}^{L}{\cal L}_{j}\left(h,r_{j},x,\widehat{y}_{<j},y\right)$ where $r=\begin{bmatrix}r_{1}&\dots&r_{L}\end{bmatrix}^{T}$. The ${\cal L}^{\text{Token}}$-risk is denoted by ${\cal R}_{\text{Token}}(h,r)=\mathop{\mathbb{E}}_{(x,y)\sim{\cal P}_{X,Y}}[{\cal L}^{\text{Token}}(h,r,x,y)]$.
For analysis, we assume teacher-forced training (Williams and Zipser, 1989; Lamb et al., 2016): at $j^{\text{th}}$ step, $\mathcal{L}_{j}$ is evaluated on the output of $r_{j}$, but we roll out the leftward context for step j+1j{+}1 using the oracle rejection decision $\bm{1}_{l(y,\widehat{y}^{h}_{j})\geq c_{j}(x,\widehat{y}_{<j},y)}$. This decouples the training distribution of $\widehat{y}_{<j}$ from the router’s predictions; the context depends only on $(x,y)$ via the predictor and expert outputs. In experiments, we don’t limit ourselves to teacher forcing training due to the risk of exposure bias; see Appendix H.